OpenStudy (anonymous):

How do I solve the following simultaneous equation: y = x^2 and y = 5x - 6

OpenStudy (anonymous):

x2-2x-3x+6=0 x(x-2)-3(x-2)=0 (x-2)(x-3)=0

OpenStudy (anonymous):

How did you work that out?

OpenStudy (anonymous):

Do you know about factorization ??

OpenStudy (anonymous):

Yes, but all we were told to use is the substitution method

OpenStudy (anonymous):

Here firstly we use Substitution Method only.. See you have : \[y = x^2\]

OpenStudy (anonymous):

In second equation: \[y = 5x - 6\] Here put that value of y.. \[x^2 = 5x - 6\] Getting ??

OpenStudy (anonymous):

Oh yes